# biblio.bib Resolution Limiting Factors in Low-Energy Cascade Zenith Reconstruction with the IceCube Upgrade

### v1

@@ Abstract
The IceCube Neutrino Observatory includes low energy extensions such as the existing DeepCore subarray and the upcoming IceCube Upgrade, which will consist of seven new strings of photosensors with denser instrumentation than the existing array. The setup will allow for the study of neutrino oscillations with greater sensitivity compared to the existing instrumentation, improve neutrino mass ordering studies, and test for the unitarity of the PMNS mixing matrix with high precision. A critical component in these low-energy physics analyses is the accurate reconstruction of event information, particularly the zenith angle of incoming neutrinos. In this study, we discuss the processes that limit the zenith resolution, which include the transverse spread of the hadronic shower, in-ice photon scattering, module resolutions, and module noise. By considering approximations to these processes, we aim to approach the intrinsic zenith resolution limits for purely hadronic events.
@@ 1 Introduction
The IceCube Neutrino Observatory consists of a cubic kilometer of ice instrumented with 5160 Digital Optical Modules (DOMs) distributed across 86 strings, positioned at depths ranging from 1450 m to 2450 m. Most of the detector features modules spaced 125 m apart horizontally and 17 m vertically, optimized for detecting astrophysical neutrinos with energies exceeding 100 GeV [Aartsen:2017]. The bottom center of the detector includes 8 strings forming the DeepCore region where the optical transparency of the ice layers is the cleanest in the instrumented region [Abbasi:2012]. Through its increased instrumentation density, DeepCore lowers the detection energy threshold to approximately 10 GeV [Abbasi:2012] enabling improved event reconstruction of such low-energy events and allowing for the exploration of the fundamental properties of neutrinos [Osc_analysis]] [numu_appearance] [NSI_analysis] [Sterile_analysis].
@@ 1.1 The IceCube Upgrade
The IceCube Upgrade [ishihara2019icecube] will be deployed in the 2025/2026 Antarctic Field season, where seven additional strings of photosensors and calibration devices will be dispersed within the DeepCore volume. Most of the new optical sensors will be installed at depths between 2160 m and 2430 m, with a smaller number of calibration devices placed between 1300 m and 1900 m as well as below 2450 m. The modules will be positioned with an average horizontal and vertical spacing of 30 m and 3 m, respectively, designed to further lower the energy threshold to approximately 3 GeV [eller2023sensitivity]. The increased instrumentation density will enhance photon statistics on a per-event basis and consequently improve event reconstruction at a few GeVs. The majority of the new optical modules will be multi-PMT Digital Optical Modules (mDOMs) [mDOM], each equipped with 24 3" PMTs, along with several dual-PMT modules called DEggs111Dual optical sensors in an Ellipsoid Glass for Gen2 [DEgg]. Compared to the currently deployed single-PMT modules and DEggs, the mDOMs offer enhanced directional coverage and nearly isotropic sensitivity allowing for more accurate reconstruction of neutrino direction.
@@ 1.2 Event signatures
Neutrinos are detected by underwater/ice neutrino telescopes when they interact with the transparent detector medium, producing relativistic charged particles that emit Cherenkov radiation detectable by the optical modules [LU2020163678]. These events are classified as "cascades" or "tracks" depending on the light signature left in the detector by the produced charged particles. Above a few GeVs, neutrinonucleon interactions are typically deep inelastic scatterings [Bustamante:2019] which result in the disintegration of the nucleon and a subsequent hadronic shower, observable in all events starting inside the detector volume. The emitted Cherenkov photons undergo scattering as they propagate through the medium toward the modules which distorts their initial directional information. The finite number of PMTs in the receiving modules results in a finite photon directional resolution, which improves proportionally with the PMT count per module. Noise in the IceCube modules is noncorrelated Poissonian distributed noise of a thermal origin and correlated noise originating from scintillation within the module glass [Larson_thesis]. These generate signals mimicking physics hits, further complicating the extraction of actual physics information.
@@
In CC and flavorindependent neutralcurrent (NC) interactions, electromagnetic and hadronic showers are observed while charged-current (CC) interactions produce an additional muon which leaves a track signature. In CC interactions, 17% of the tau leptons decay to muons [ParticleDataGroup:2020ssz] while the rest decay to cascades. At energies around 10 GeV, due to the coarse detector granularities, distinguishing between these different topologies becomes increasingly challenging [osti_10349667]. In CC events, the track lever arm can be short enough that it becomes difficult to distinguish from a cascade event while in interactions, the small spatial separation between the dual cascades causes them to resemble a single cascade. These events can thus be modeled well by a point-like singlecascade model in reconstructions which are described by relatively fewer parameters than tracks. Given the similarity of event topologies at these very low energies and motivated by the simpler event modeling for single cascades, this study focuses on the directional reconstruction for NC events characterized by a hadronic shower.
1.3 Importance of zenith reconstruction
@@
This study aims to analyze the processes that limit the zenith directional information in IceCube Upgrade events. This is directly relevant for all low-energy neutrino oscillation analyses in water (ice) Cherenkov experiments [Osc_analysis] since it translates into the oscillation length , given by (where R denotes the radius of the Earth), as depicted in Fig. 1. Previous studies in IceCube have shown that a 5%, 10%, and 15% improvement in the zenith angle resolution results in a 3%, 10%, and 15% enhancement in Neutrino Mass Ordering (NMO) sensitivity, respectively [Jan_thesis]. Since better reconstruction yields improvements in analysis sensitivities, efforts are directed toward improving reconstruction algorithms. Therefore, it becomes essential to check the performance status of the reconstruction algorithms relative to the resolution limits, the latter of which are determined by assuming perfect detector modeling and accounting for all processes that constrain the achievable resolutions.
@@
The study is motivated by a prior investigation pursued by the KM3NeT Collaboration222The Cubic Kilometre Neutrino Telescope (KM3NeT) [KM3NeT:2017], which explored the information-limiting processes in CC and CC events in the ORCA333Oscillation Research with Cosmics in the Abyss (ORCA) water Cherenkov detector. These included fluctuations in the hadronic shower composition, particle propagation, and photon sampling effects. The resolution limiting processes included in this study are the transverse shower spread, in-ice light scattering, module directional resolutions due to finite PMTs in the modules, and noise in the PMTs (Section 5).
@@ 2 Approximations in this study
Since events in Cherenkov detectors are characterized by detected photon signatures, the reconstruction of events relies on the precise modeling of the medium through which the photons propagate. In the IceCube experiment, the optical properties of glacial ice show a strong depth dependence which reflects the changes in Earth’s climate history [Aartsen:2013]. The ice sheet is also not horizontally uniform and is titled due to the underlying bedrock topography [chirkin2023improved]. In addition, there is a direction-dependent anisotropy effect due to birefringent polycrystals which leads to an average deflection in the diffusion pattern towards the flow axis [IceCube_ice]. Moreover, the modules are deployed into holes made by hot water drilling. This hole ice [fiedlschuster2019effect], on re-freezing, has different optical properties than the bulk ice and contains a bubbly column (with poor optical properties) approximately at the center of the column.
Due to these ice anisotropies and the grid-like geometry of the detector, the measured angular and timing distribution of photons (Fig. 4) by a receiving optical module depends strongly on the interaction vertex position and event direction. The similarity of the scattering length and module spacing scales further make the photon distributions analytically not tractable. In an ideal likelihood-based reconstruction, the actual expected
@@
and direction of the event. Given the detector’s large volume and the extensive number of optical
@@
infeasible. In practical application, simplified photon distribution models using approximate symmetries and translational invariance are employed. In this study, in contrast, we circumvent the modeling challenges by simplifying the assumptions about the photon propagation and using other approximations in the simulation of the detector.
Events are simulated and reconstructed in a homogeneous ice medium, where the optical properties remain uniform throughout the volume, incorporating translational and rotational symmetry in the photon distributions in space and time. There is no ice stratification, layer undulations, birefringence, or hole ice, which otherwise require detailed parameterization of the ice-depth dependence. All PMTs in the modules are assumed to be identical, with the same angular acceptance, which defines the fraction of photons detected at a given angle relative to the PMT axis. Moreover, the simulations in this study only include 24-PMT mDOMs at all module positions in the detector geometry whereas typical MC simulations also include other module types such as pDOMs and DEggs. This results in better directional resolutions since mDOMs offer a better pointing resolution as compared with other modules with fewer PMTs. The assumption is appropriate for this study which seeks to estimate the optimistic resolution limits.
@@
To generate the expected angular and timing distribution of detected photons relative to the shower axis and vertex time (Section 4.2), respectively, a randomized geometry (Section 4.3) is created. Events are simulated with interaction vertices distributed uniformly throughout this geometry. The vertex-averaged distribution of detected photons across all modules removes photon sampling bias caused by the local geometry around the event. This approach allows the use of a single probability density function (PDF), representing the modeled angular and timing distribution of photons summed over all hit modules and averaged over multiple events, to reconstruct all events. The PDFs generated in this way are independent of the source-receiver distance and represent an idealization, deviating from standard likelihood reconstructions (Section 4.1) that rely on photon distributions from each module. The expected number of photon hits in a module, defined as charge in this study, is distance-dependent and has been incorporated in the likelihood formulation. Note that the definition of charge in this study slightly differs from the one used by the IceCube Collaboration [SPE_template], as explained later in this section.
@@ 3 Event Simulation
This study uses the standalone, open-source version of the Photon Propagation Code (PPC) [PPC] used by the IceCube Collaboration. Events are simulated as purely hadronic cascades resulting from NC interactions mediated by a neutral boson, with no charged lepton emitted at the interaction vertex. The standard input to PPC is comprised of information on an IceCube event which includes the event type, interaction vertex coordinates and time, incoming directions in zenith and azimuth, and the energy of the hadronic shower. Since the study focuses on intrinsic directional resolution limits, the uncertainties linked with angular offset between the parent neutrino and the shower axis, have not been considered. However, it is worth mentioning that the average kinematic opening angle between the incoming neutrino and the shower axis scales with energy [kinematic_angle:2022] approximately as . The coefficients = 12∘ and = 0.3 were fitted to the angular distribution of shower secondaries in 10 GeV purely hadronic cascades using the GENIE [Andreopoulos:2010] event generator.
@@
The Cherenkov photons emitted from propagating charged particles in the shower are simulated using light yield parameterizations [Wiebusch:1995bw] of physical processes and then propagated through the ice until they get absorbed or reach one of the modules. The scattering angle of propagating photons at each scattering point is sampled from a linear combination of scattering functions [Aartsen:2013], which is predominantly characterized by forward scattering. The output from PPC includes the strings and modules hit by photons with the corresponding timestamps, photon directions at the point of incidence on a module, and the polar coordinates of the impact point relative to the DOM center.
@@
The overall angular distribution of emitted photons by the individual charged particles in the hadronic shower relative to the shower axis, referred to as "shower spread" in this study, is given as , where is the length of the light emitting track segment and = 1-cos() [Wiebusch:1995bw]. To switch off the shower spread, the simulation skips the step in the simulation that samples the opening angles of emitted photons from the above-described distribution and instead draws them from a Cherenkov cone, whose axis aligns with the shower axis, with a wavelength-dependent opening angle [PhysRev.52.378]. The scattering and absorption of photons are governed by the bulk ice properties and have been derived from [Aartsen:2013]. For this study, the scattering and absorption coefficients are set to and , respectively. These values represent the ice properties at a depth of approximately 2300 m, corresponding to the most transparent ice layers within the Upgrade physics region.
@@
The PPC simulations only involve the generation and propagation of photons through the ice. Other information loss factors such as module resolutions and module noise for this study are manually incorporated to the PPC output. To simulate a multi-PMT Digital Optical Module (mDOM) [mDOM], which will be the most common optical module in the Upgrade, 24 PMTs are arranged in four belts positioned at zenith angles of , , , and relative to the center of the module. The PMT population per belt is 5, 7, 7, and 5, respectively, with the PMTs within each belt uniformly distributed in azimuth. The overall angular resolution of the optical modules, referred to as "module resolution" in this study, is simulated using the following procedure to account for the finite number of PMTs and the shape of the angular acceptance curve for each PMT.
@@
The PMT closest to the point of photon impact is first chosen as shown in Fig. 2. The angle between the incident photon direction and the PMT axis vector () is then calculated. Rejection sampling is employed to select photons based on this angle and the angular acceptance curve of a PMT. For the purpose of our reconstruction, the PMT axis (P, P) is assumed as the photon direction for the sampled photons reflecting the known polar orientation of the hit PMT. The procedure is illustrated in Fig. 2 using two photons, and , incident on a multi-PMT module. The angle between and the nearest PMT vector is smaller than the angle between and . Consequently, has a higher probability of registering a photon hit as the acceptance function (shown in orange) increases monotonically with decreasing opening angles. The simulation method here does not preserve the absolute detection efficiency [Abbasi:2010] of the mDOMs. However, this detail is inconsequential for our study because all resolutions are binned in photon hit counts. Whenever module resolutions are not included in the simulation, the true direction of the photon is used instead.
@@
Signal purity in an IceCube event is defined as the fraction of surviving physics hits after noise-cleaning out of the total observed photon hits, which includes correlated and uncorrelated noise hits. The overall dark rate for the mDOM PMTs was measured to be 750Hz [mDOM_dark_rate] which translates to 0.1 noise hits per module in a waveform readout window of approximately 6s. As a first-order approximation, the number of noise hits for a module is assumed to be Poisson distributed with =0.1. This corresponds to a high signal purity of approximately 90% which is in agreement with the values given in [eller2023sensitivity]. To simulate a noise hit in a module, a PMT is randomly selected and a timestamp is sampled uniformly within the event time window. The overall noise rate is included in the simulation without distinguishing between correlated and uncorrelated noise and therefore, without any spatial or temporal clustering of noise hits. The generated noise hits are then combined with the physics hits to give the total observed hits for an event. Given the negligible impact of noise on our reconstruction (Section 5) due to the high signal purity, this technique provides a sufficiently good approximation of noise contamination in IceCube events.

### v2

# biblio.bib Resolution Limiting Factors in Low-Energy Cascade Zenith Angle Reconstruction with the IceCube Upgrade

@@ Abstract
The IceCube Neutrino Observatory includes low energy extensions such as the existing DeepCore subarray and the upcoming IceCube Upgrade, which will consist of seven new strings of photosensors with denser instrumentation than the existing array. The setup will allow for the study of neutrino oscillations with greater sensitivity compared to the existing instrumentation, improve neutrino mass ordering studies, and test for the unitarity of the PMNS mixing matrix with high precision. A critical component in these low-energy physics analyses is the accurate reconstruction of event information, particularly the zenith angle of incoming neutrinos. In this study, we discuss the processes that limit the zenith angle resolution, which include the transverse spread of the hadronic shower, in-ice photon scattering, module resolutions, and module noise. By considering approximations to these processes we aim to approach the intrinsic zenith angle resolution limits for neutral-current events characterized by a hadronic shower.
@@ 1 Introduction
The IceCube Neutrino Observatory consists of a cubic kilometer of ice instrumented with 5160 Digital Optical Modules (DOMs) distributed across 86 strings, positioned at depths ranging from 1450 m to 2450 m. Most of the detector features modules spaced 125 m apart horizontally and 17 m vertically, optimized for detecting astrophysical neutrinos with energies exceeding 100 GeV [Aartsen:2013]. The bottom center of the detector includes 8 strings forming the DeepCore region where the optical transparency of the ice layers is the cleanest in the instrumented region [Abbasi:2012]. Through its increased instrumentation density, DeepCore lowers the detection energy threshold to approximately 10 GeV [Abbasi:2012] enabling improved event reconstruction of such low-energy events and allowing for the exploration of the fundamental properties of neutrinos [Osc_analysis, numu_appearance, NSI_analysis, Sterile_analysis].
@@ 1.1 The IceCube Upgrade
The IceCube Upgrade [ishihara2019icecube] will be deployed in the 2025/2026 Antarctic Field season, where seven additional strings of photosensors and calibration devices will be dispersed within the DeepCore volume. Most of the new optical sensors will be installed at depths between 2160 m and 2430 m, with a smaller number of calibration devices placed between 1300 m and 1900 m as well as below 2450 m. The modules will be positioned with an average horizontal and vertical spacing of 30 m and 3 m, respectively, designed to further lower the energy threshold to approximately 3 GeV [eller2023sensitivity]. The increased instrumentation density will enhance photon statistics on a per-event basis and consequently improve event reconstruction at a few GeV. The majority of the new optical modules will be multi-PMT Digital Optical Modules (mDOMs) [mDOM], each equipped with 24 3" PMTs, along with several dual-PMT modules called DEggs111Dual optical sensors in an Ellipsoid Glass for Gen2 [DEgg]. Compared to the currently deployed single-PMT modules and DEggs, the mDOMs offer enhanced directional coverage and nearly isotropic sensitivity allowing for more accurate reconstruction of neutrino directions.
@@ 1.2 Event signatures
Neutrinos are detected by underwater or ice-based neutrino telescopes when they interact with the transparent detector medium or surrounding rock, producing relativistic charged particles. As these particles traverse the detector, they emit Cherenkov radiation which is recorded by the optical modules deployed within the detector [LU2020163678]. The events are classified as "cascades" or "tracks" depending on the light signature left in the detector by the produced charged particles. Above a few GeVs, neutrino-nucleon interactions are typically deep inelastic scatterings [Bustamante:2019] which result in the disintegration of the nucleon and a subsequent hadronic shower, observable in all events starting inside the detector volume. The emitted Cherenkov photons undergo scattering as they propagate through the medium toward the modules which distorts their initial directional information. The finite number of PMTs in the receiving modules results in a finite photon directional resolution, which improves proportionally with the PMT count per module. Noise in the IceCube modules is non-correlated Poissonian distributed noise of a thermal origin and correlated noise originating from scintillation light due to natural radioactive decays within the module glass [Larson_thesis]. These generate signals mimicking physics hits, further complicating the extraction of actual physics information.
@@
In CC and flavor-independent neutral-current (NC) interactions, electromagnetic and hadronic showers are created while charged-current (CC) interactions produce an additional muon which leaves a track signature. In CC interactions, a hadronic shower is formed at the interaction vertex, followed by either a hadronic or leptonic shower from the subsequent tau decay, with known branching ratios [ParticleDataGroup:2020ssz]. At energies around 10 GeV, due to the coarse detector granularities, distinguishing between these different topologies becomes increasingly challenging [osti_10349667]. In CC, CC and NC events, there are no experimentally observed differences due to negligible cascade elongation [Steuer_theses] and tau decay [nu_tau] lengths compared to the inter-string spacing. All of these events can thus be modeled well by a point-like single-cascade model in reconstructions which are described by relatively fewer parameters than muon tracks. Given the similarity of event topologies at these very low energies and motivated by the simpler event modeling, this study focuses on the directional reconstruction for NC events characterized by a hadronic shower.
1.3 Importance of zenith angle reconstruction
@@
This study aims to analyze the processes that limit the zenith angle directional information in IceCube Upgrade events. This is directly relevant for all low-energy neutrino oscillation analyses in water (ice) Cherenkov experiments [Osc_analysis] since it translates into the oscillation length , given by (where R denotes the radius of the Earth and is the zenith angle of the incoming neutrinos), as depicted in Fig. 1. Previous studies in IceCube have shown that a 5%, 10%, and 15% improvement in the zenith angle resolution results in a 3%, 10%, and 15% enhancement in Neutrino Mass Ordering (NMO) sensitivity, respectively [Jan_thesis]. Since better reconstruction yields improvements in analysis sensitivities, efforts are directed toward improving reconstruction algorithms. It is therefore essential to check the performance status of the reconstruction algorithms relative to the resolution limits, which are determined under the assumption of ideal detector modeling and by incorporating all physical processes that constrain the achievable resolutions.
@@
The study is motivated by a prior investigation pursued by the KM3NeT Collaboration222The Cubic Kilometre Neutrino Telescope (KM3NeT) [KM3NeT:2017], which explored the information-limiting processes in CC and CC events in the ORCA333Oscillation Research with Cosmics in the Abyss (ORCA) water Cherenkov detector. These included fluctuations in the hadronic shower composition, particle propagation, and photon sampling effects. The resolution limiting processes included in this study are the transverse shower spread, in-ice light scattering, module directional resolutions due to finite PMTs in the modules, and noise in the PMTs (Section 5). Given the lack of topological distinction between NC and CC cascade events at these low energies, the resolution limits derived in this study are equally applicable to CC and CC interactions, with only minimal corrections. The CC events show a strong zenith angle dependence because of their flavor specific nature and the baseline-dependent neutrino flavor oscillations. As a result, they contribute significantly to IceCube oscillation analyses [Osc_2023, Osc_2018, Osc_analysis].
@@ 2 Approximations in this study
Since events in Cherenkov detectors are characterized by detected photon signatures, the reconstruction of events relies on the precise modeling of the medium through which the photons propagate. In the IceCube experiment, the optical properties of glacial ice show a strong depth dependence which reflects the changes in Earth’s climate history [Aartsen:2013]. The ice sheet is also not horizontally uniform and is titled due to the underlying bedrock topography [chirkin2023improved]. In addition, there is a direction-dependent anisotropy effect due to birefringent polycrystals which leads to an average deflection in the diffusion pattern towards the flow axis [IceCube_ice]. Moreover, the modules are deployed into holes made by hot water drilling. This hole ice [fiedlschuster2019effect], on re-freezing, has different optical properties than the bulk ice and contains a bubbly column (with poor optical properties) approximately in the center of the column.
Due to these ice anisotropies and the grid-like geometry of the detector, the measured angular and timing distribution of photons (Fig. 4) by a receiving optical module depends strongly on the interaction vertex position and event direction. The similarity of the scattering length and module spacing scales further make the photon distributions analytically intractable. In an ideal likelihood-based reconstruction, the actual expected
@@
and direction of the event. Given the detector’s extensive volume and the large number of optical
@@
infeasible. In practical application, simplified photon distribution models using approximate symmetries and translational invariance are employed [improvedcascadereco:2024csv]. In this study, in contrast, we circumvent the modeling challenges by simplifying the assumptions about the photon propagation and using other approximations in the simulation of the detector.
Events are simulated and reconstructed in a homogeneous ice medium, where the optical properties remain uniform throughout the volume, incorporating translational and rotational symmetry in the photon distributions in space and time. There is no ice stratification, layer undulations, birefringence, or hole ice, which otherwise require detailed parameterization of the ice-depth dependence. All PMTs in the modules are assumed to be identical, with the same angular acceptance, which defines the fraction of photons detected at a given angle relative to the PMT axis. Moreover, the simulations in this study only include 24-PMT mDOMs at all module positions in the detector geometry whereas typical MC simulations also include other module types such as pDOMs and DEggs. This results in better directional resolutions since mDOMs offer a better pointing resolution as compared with other modules with fewer PMTs. The assumption is appropriate for this study which seeks to estimate optimistic resolution limits.
@@
To generate the expected angular and timing distribution of detected photons relative to the shower axis and vertex time (Section 4.2), respectively, a randomized geometry (Section 4.3) is created. Events are simulated with interaction vertices distributed uniformly throughout this geometry. The vertex-averaged distribution of detected photons across all modules removes the photon sampling bias caused by the local geometry around the event. This approach allows the use of a single probability density function (PDF), representing the modeled angular and timing distribution of photons summed over all hit modules and averaged over multiple events, to reconstruct all events. The PDFs generated in this way are independent of the source-receiver distance and represent an idealization, deviating from standard likelihood reconstructions (Section 4.1) that rely on photon distributions from each module. The expected number of photon hits in a module, defined as charge in this study, is distance-dependent and has been incorporated in the likelihood formulation. Note that the definition of charge in this study differs slightly from the one used by the IceCube Collaboration [SPE_template], as explained later in this section.
@@ 3 Event Simulation
This study uses the standalone, open-source version of the Photon Propagation Code (PPC) [PPC] used by the IceCube Collaboration. Events are simulated as purely hadronic cascades resulting from NC interactions mediated by a neutral boson, with no charged lepton emitted at the interaction vertex. The standard input to PPC is comprised of information on an IceCube event which includes the event type, interaction vertex coordinates and time, incoming directions in zenith and azimuth, and the energy of the hadronic shower. Since only the directions of outgoing charged particles within the detector are experimentally observable, the study focuses exclusively on the intrinsic directional resolution limits. The uncertainties associated with the kinematic opening angle between the incoming neutrino and shower axis are not included. The average kinematic angle between the incoming neutrino and the shower axis varies with energy [kinematic_angle:2022] approximately as . The coefficients = 12∘ and = 0.3 were fitted to the angular distribution of shower secondaries in 10 GeV purely hadronic cascades using the GENIE [Andreopoulos:2010] event generator.
@@
The charged particles in the shower are generated and propagated using GEANT3 simulation software [Geant3] (version 3.16). The hadronic interactions within ice are modeled with the GHEISHA 7.0 package [GHEISHA] integrated within the GEANT3 framework. The Cherenkov photon yield is determined based on the energy loss of the charged particles, and the parameterisations [Wiebusch:1995bw] are directly used by PPC to determine the distribution of Cherenkov photons. The photons are then propagated by PPC through the ice until they get absorbed or reach one of the modules. The scattering angle of propagating photons at each scattering point is sampled from a linear combination of scattering functions [Aartsen:2013], which is predominantly characterized by forward scattering. The output from PPC includes the strings and modules hit by photons with the corresponding timestamps, photon directions at the point of incidence on a module, and the polar coordinates of the impact point relative to the DOM center.
@@
The overall angular distribution of photons emitted by the individual charged particles in the hadronic shower relative to the shower axis, referred to as "shower spread" in this study, is given as , where represents the length of the light-emitting track segment. The variable is defined as = 1-cos() [Wiebusch:1995bw], where denotes the angle between the emitting charged particle track and emitted photon direction. To switch off the shower spread, the simulation skips the step in the simulation that samples the emission angles of photons from the above-described distribution and instead draws them from a Cherenkov cone, whose axis aligns with the shower axis, with a wavelength-dependent cone angle [PhysRev.52.378]. The scattering and absorption of photons are governed by the bulk ice properties and have been derived from [Aartsen:2013]. For this study, the scattering and absorption coefficients are set to and , respectively. These values represent the ice properties at a depth of approximately 2300 m, corresponding to the most transparent ice layers within the Upgrade physics region.
@@
The PPC simulations only involve the generation and propagation of photons through the ice. The PPC output is adjusted to include other information loss factors such as module resolutions and module noise. To simulate a multi-PMT Digital Optical Module (mDOM) [mDOM], which will be the most common optical module in the Upgrade, 24 PMTs are arranged in four belts positioned at zenith angles of , , , and relative to the center of the module. The PMT population per belt is 5, 7, 7, and 5, respectively, with the PMTs within each belt uniformly distributed in azimuth. The overall angular resolution of the optical modules, referred to as "module resolution" in this study, is simulated using the following procedure to account for the finite number of PMTs and the shape of the angular acceptance curve for each PMT.
@@
The PMT closest to the point of photon impact is first chosen as shown in Fig. 2. The angle between the incident photon directions and the PMT axis vectors, such as or , is then calculated. Rejection sampling is employed to select photons based on this angle and the angular acceptance curve of a PMT. For our reconstruction, PMT axis vectors, such as or , are assumed as the photon directions for the sampled photons reflecting the known polar orientation of the hit PMTs. The procedure is illustrated in Fig. 2 using two photons, and , incident on a multi-PMT module. The angle between and the nearest PMT vector is smaller than the angle between and . Consequently, has a higher probability of registering a photon hit as the acceptance function (shown in orange) increases monotonically with decreasing angles. The simulation method here does not preserve the absolute detection efficiency [Abbasi:2010] of the mDOMs. However, this detail is inconsequential for our study because all resolutions are shown as a function of photon hit counts. Whenever module resolutions are not included in the simulation, the true direction of the photon is used instead.
@@
The signal purity of an IceCube event is defined as the fraction of the total observed photon hits that are physics hits, after noise cleaning. The total observed photon hits also include noise hits which are classified as correlated noise hits and uncorrelated noise hits. The overall dark rate of each PMT in mDOMs was measured to be 750Hz [mDOM_dark_rate] which translates to 0.1 noise hits per module in a waveform readout window of approximately 6s. As a first-order approximation, the number of noise hits for a module is assumed to be Poisson distributed with =0.1. This corresponds to a high signal purity of approximately 90% which is in agreement with the values given in [eller2023sensitivity]. To simulate a noise hit in a module, a PMT is randomly selected and a timestamp is sampled uniformly within the event time window. The overall noise rate is included in the simulation without distinguishing between correlated and uncorrelated noise and therefore, without any spatial or temporal clustering of noise hits. The generated noise hits are then combined with the physics hits to give the total observed hits for an event. Given the negligible impact of noise on our reconstruction (Section 5) due to the high signal purity, this technique provides a sufficiently good approximation of noise contamination in IceCube events.